Bar Delivery Company purchased a new delivery truck for $65,400 on April 1, 2019. The truck is expected to have a service life of 5 years or 170,400 miles and a residual value of $3,120. The truck was driven 9,500 miles in 2019 and 11,900 miles in 2020. Bar computes depreciation expense to the nearest whole month.

Required: Compute depreciation expense for 2019 and 2020 using the following methods: (Round your answers to the nearest dollar.)

Straight-line method 2019 $ 2020 $ Sum-of-the-years'-digits method 2019 $ 2020 $ Double-declining-balance method 2019 $ 2020 $ Activity method 2019 $ 2020 $ For each method, what is the book value of the machine at the end of 2019? At the end of 2020? (Round your answers to the nearest dollar.) Straight-line method 2019 $ 2020 $ Sum-of-the-years'-digits method 2019 $ 2020 $ Double-declining-balance method 2019 $ 2020 $ Activity method 2019 $ 2020 $
Answer #1

Straight line method

Depreciation = (cost of asset - salvage value)/useful life of asset

Depreciation in Year 2019 = (65,400 - 3,120)/5 = $12,456

Depreciation in Year 2020 = $12,456

Book value at the end of 2019 = 65,400 - 12,456 = $52,944

Book value at the end of 2020 = 52,944 - 12,456 = $40,488

Sum of the year digit method

Year Depreciation Base Remaining life of asset Depreciation Fraction Dep Exp Book value
2019 65,400 - 3,120 = 62,280 5 5/15 20,760 44,640
2020 62,280 4 4/15 16,608 28,032

1 + 2+ 3+ 4+ 5 = 15

Double declining Method

Depreciation rate = (1/useful life)*100*2 = (1/5)*100*2 = 40%

Depreciation in Year 2019 = cost of asset * rate of dep = 65,400*40% = $26,160

Depreciation in Year 2020 = (65,400 - 26,160)*40% = $15,696

Book value at the end of 2019 = 65,400 - 26,160 = $39,240

Book value at the end of 2020 = 39,240 - 15,956 = $23,284

Activity Method

Depreciation = (cost of asset - salvage value)* period activity/ Lifetime activity

Depreciation in Year 2019 = (65,400 - 3,120)*9,500/170,400 = $3,472

Depreciation in Year 2020 =(65,400 - 3,120)*11,900/170,400 = $4,349

Book value at the end of 2019 = 65,400 - 3,472 = $61,928

Book value at the end of 2020 = 61,928 - 4,349= $57,579